We have a jar of coins, all pennies and quarters. All together, we have 191 coins, and the total value of all coins in the jar is $24.23. How many pennies are there in the jar?
Answer by mananth(16946)

We have a jar of coins, all pennies and quarters. All together, we have 191 coins, and the total value of all coins in the jar is $24.23. How many pennies are there in the jar?
$24.23 = 2423 cents
Let there be p pennies and q quarters
coin count equation
p+q=191...............................................(1)
Money value equation
p+25q= 2423 ......................................(2)

subtract (1) from (2)

24 q= 2232
q = 2232/24
q= 93 quarters
p = 191-93= 98 pennies



98+93*25= 2423